40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Wed Oct 20 11:31:06 2021, shruti, Update, PSOMA, Towards locking PSOMA with cantilever 
    Reply  Tue Oct 26 10:01:54 2021, shruti, Update, PSOMA, Towards locking PSOMA with cantilever TiltedCantilever.pdfcantileversurface.pdfclampsurface.pdf
       Reply  Thu Nov 18 11:09:24 2021, shruti, Update, PSOMA, Towards locking PSOMA with cantilever CavitySetup.pdfGeometry01.jpeg
          Reply  Wed Dec 1 13:21:54 2021, shruti, Update, PSOMA, Towards locking PSOMA with cantilever IMG_1250.pngIMG_1252.pngIMG_1251.pngIMG_1253.pngIMG_1254.png
             Reply  Thu Dec 2 13:40:57 2021, shruti, Update, PSOMA, Progress and cavity parameters/ lock parameters 
Message ID: 2828     Entry time: Thu Nov 18 11:09:24 2021     In reply to: 2825     Reply to this: 2829
Author: shruti 
Type: Update 
Category: PSOMA 
Subject: Towards locking PSOMA with cantilever 

Since clamping the cantilever, I began to align the triangular cavity with the cantilever.

Attachment 1 shows the setup of the cavity with the input coupler at M1, curved mirror at M2, and a cantilever with a flat optic at M3. The cantilever does not possess any knobs for fine adjustment and its initial mounting is fixed. M1 and M2 can be adjusted, but to get to TEM00 resonance, and any fine alignment, the two steering mirrors before the cavity were used.

Weird things happen while aligning:

1. While aligning, I began to see resonant peaks corresponding to the cantilever's mechanical resonance showing that the cantilever has a very high Q. I mistakenly thought the cavity was close to being aligned to resonance in this configuration while also being very stable, but later understood that the strong oscillations actually seemed more like accidental misalignments creating an optical lever. The green traces in the videos are the measured transmitted light through the cantilever. When the cavity is really aligned close to the TEM00 mode the transmission looks like the green trace here, and the reflection (from M1) is the blue trace.

2. Another time I thought it was aligned I was looking at the camera at the transmission of the cantilever [M3 in Attachment 1] and could see a single flashing spot but when I moved the camera around I actually ended up seeing two beams instead, simultaneously flashing. Minor changes in the alignment resulted in the beams individually and simultaneously turning into higher order modes. Initially I thought that the situation was as depicted in Attachment 2 and found that it was a consistent geometric solution, but later Koji pointed out that was the geometry that creates the TEM01 mode. In Attachment 2, the initial beam (red, M1->M2) overlaps not with the beam in that path from the second round trip (yellow, M1->M2) but with the one from the third round trip. Was it possible that this misalignment was too severe that instead of being TEM01 two separate, but identical, resonant cavities were created? A measurement of the FSR, which I did not do, would have proved that that was the case.

Since the transfer of fiber components into the box, I have re-aligned the cavity to the TEM00 mode. Note: replacing the fiber at the fiber launch and also changing the polarization at the fiber launch moves the beam around.

PDH locking stuff:

[more information and data coming soon]

1. The PDH signal seems too low, possibly because the LO and RF relative phase must be re-adjusted.

Attachment 1: CavitySetup.pdf  1.607 MB  Uploaded Mon Nov 22 09:55:05 2021  | Hide | Hide all | Show all
CavitySetup.pdf
Attachment 2: Geometry01.jpeg  542 kB  Uploaded Mon Nov 22 10:21:59 2021  | Show | Hide all | Show all
ELOG V3.1.3-